OpenStudy (anonymous):

If y varies inversely as the cube root of x, and y = 12 when x = 8, find y when x = 1.

OpenStudy (anonymous):

y = k/x^(1/3) 12 = k/8^(1/3) = k/2 so k = 24 y = 24/x^(1/3) y = 24/1^(1/3) = 24/1 = 24

OpenStudy (anonymous):

so my final answer is y=24?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I have another one but i will put it in another question